if 7g of water is at 10oC, and a copper block at 113oC is added to it, and the final temperature of the mixture is 54oC, what is the mass of the copper?
1
Expert's answer
2020-12-06T13:54:56-0500

| Q(H2O) | = | Q(Cu) |

Tend (H2O) = Tend (Сu) = 54 0C

c(H2O)*m(H2O)*T(H2O) = c(Cu)*m(Cu)*T(Cu)


4.18J/[g*K] * 7g * |10-54|0C = 0.385J/[g*K] * m(Cu) * |113-54|0C

1287.44 = 22.7m(Cu)

m(Cu) = 56.7 g